can you help me with this fuction plz how do you work out x of this fuction machien here it is
x-x2-+12=512

Answer by checkley71(8403)

IF YOU HAVE THE RIGHT EQUATION AS FOLLOWS --
X-X^2-12=512 THEN
-X^2-X-514=0 REMOVING THE - SIGNS BY MULTIPLYING ALL TERMS BY -1 WE GET
X^2+X+514=0 NO SIMPLE FACTORING SO WE NEED THE QUADRATIC FORMULA
X=[-B+-SQRT(B^2-4AC)]/2A
X=[-1+-SQRT(1-4*1*514)]/2
X=[-1+-SQRT(1-2056)]/2
X=[-1+-SQRT(-2055)]/2
X=-1/2+45.332i/2
X=-.5+22.666i
AND
X=-1/2-45.332i/2
X=-.5-22.666i